A simple random sample of size n = 40 is drawn from a population. The sample mean is found to be x = 120.2 and the sample staThe upper bound is ___
Answer #1

solution

Given that,

\bar x = 120.2

s =12.7

n = 40

Degrees of freedom = df = n - 1 =40 - 1 = 39

At 99% confidence level the t is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t\alpha /2  df = t0.005, 39=   2.708 ( using student t table)

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.708* ( 12.7/ \sqrt 40) = 5.44

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

120.2 - 5.44 < \mu <120.2 + 5.44

114.76 < \mu < 125.64

(lower bound= 114.76 , upper bound=125.64)